OpenStudy (anonymous):

What is the value of "h' in the function y=2^(x+1) ?

OpenStudy (tkhunny):

There is no "h". Please ask a better or more complete question.

OpenStudy (freethinker):

you are quite persistent knowing the value of "h" :)

OpenStudy (freethinker):

http://openstudy.com/study#/updates/51688289e4b0344769c98110 here is your previous post :P

OpenStudy (anonymous):

Thats all there giving , so thats all that im asking. Please dont act rude !

OpenStudy (anonymous):

And Lol Me ?? No my teacher is . @freethinker

jimthompson5910 (jim_thompson5910):

If you start with y = 2^x then you can translate, dilate, reflect etc using this form y = a*2^(x-h) + k where the variable a handles dilation/compression and reflection h does horizontal translation k does vertical translation

OpenStudy (anonymous):

@tkhunny SEE !

OpenStudy (anonymous):

And what might that be ? @freethinker

jimthompson5910 (jim_thompson5910):

y=2^(x+1) y=1*2^(x+1) y=1*2^(x+1) + 0 y=1*2^(x-(-1)) + 0 Compare this to y = a*2^(x-h) + k